### v1

# Enhancing Arabic Automated Essay Scoring with Synthetic Data and Error Injection

Abstract
[…]
utf8
Enhancing Arabic Automated Essay Scoring
with Synthetic Data and Error Injection
Chatrine Qwaider,1 Bashar Alhafni,1,2 Kirill Chirkunov,1 Nizar Habash,1,2 Ted Briscoe1 1MBZUAI, 2New York University Abu Dhabi {chatrine.qwaider,kirill.chirkunov,ted.briscoe}@mbzuai.ac.ae {alhafni,nizar.habash}@nyu.edu
[…]
utf8
Enhancing Arabic Automated Essay Scoring
with Synthetic Data and Error Injection
Chatrine Qwaider,1 Bashar Alhafni,1,2 Kirill Chirkunov,1 Nizar Habash,1,2 Ted Briscoe1 1MBZUAI, 2New York University Abu Dhabi {chatrine.qwaider,kirill.chirkunov,ted.briscoe}@mbzuai.ac.ae {alhafni,nizar.habash}@nyu.edu
1 Introduction
Automated Essay Scoring (AES) is a technology that automates the evaluation and scoring of essays to assess language learners’ writing quality while eliminating the need for human intervention Shermis and Burstein (2003). AES has gained great interest due to its significant benefits in the field of education Lagakis and Demetriadis (2021); Susanti et al. (2023). AES systems help teachers evaluate many essays with consistent scoring and reduced workload. On the other hand, AES helps students improve their writing quality through rapid real-time scoring and feedback Hahn et al. (2021). Unlike for English, for Modern Standard Arabic (MSA) it is difficult to develop robust and scalable AES systems, primarily due to the lack of essay datasets necessary for building effective Arabic AES Lim et al. (2021); Elhaddadi et al. (2024). This paper presents our proposed framework to tackle the issue of data scarcity and quality by utilizing Transformers and Large Language Models (LLMs) to generate and build a synthetic dataset. Our approach begins with prompting GPT-4o to generate a variety of Arabic essays covering multiple topics and different writing proficiency levels as defined by the Common European Framework of Reference (CEFR) Council of Europe (2001). Subsequently, we use a controlled error injection model to introduce errors into the correct Arabic essays, ensuring that erroneous essays reflect the type of errors that are commonly made by learners of Arabic in real-world scenarios. Our error injection approach consists of two steps: (i) Error Type Prediction, where a fine-tuned CAMeLBERT MSA model Inoue et al. (2021) classifies the most likely error type for each word, and (ii) Error Realization, where we apply a bigram MLE model to determine the most probable transformation for each predicted error type. Our framework enables the generation of realistic human-like essays, enhancing data augmentation for Arabic AES systems. Our main contributions are as follows:
  • Proposing a framework based on LLMs and Transformers for augmenting Arabic essays that accurately reflect human writing patterns.
  • Creating a synthetic Arabic AES dataset with 3,040 essays annotated with CEFR proficiency levels. [^1] The full dataset will be publicly available for research.
  • Developing an Arabic essay auto-marking system using a BERT-based scoring model, enabling accurate and scalable evaluation of Arabic essays based on CEFR standards.
The rest of the paper is organized as follows: §2 reviews related work on AES, §3 describes the dataset, and §4 outlines our data augmentation approach. §5 details the error injection methods, followed by evaluation in §6 and §7 presents conclusion and future work.
2 Related Work
The field of AES has been investigated extensively, particularly in English Lim et al. (2021); Ramesh and Sanampudi (2022), where some valuable tools have been introduced such as IntelliMetric Elliott et al. (2003), e-rater Attali and Burstein (2006), Grammarly, [^2] https://app.grammarly.com/ Write and Improve [^3] https://writeandimprove.com/ Yannakoudakis et al. (2018), and others. The development and validation of English AES have been enabled by the large-scale annotated datasets that have been released, such as the First Cambridge English (FCE) dataset [^4] https://ilexir.co.uk/datasets/index.html Yannakoudakis et al. (2011), Automated Student Assessment Prize (ASAP) dataset, [^5] https://www.kaggle.com/c/asap-aes the TOEFL11 corpus Blanchard (2013), and the ICLE (International Corpus of Learner English) Granger (2003). These datasets contained thousands of student essays with proficiency-level grades, often along multiple dimensions. In contrast, research on Arabic AES has not received comparable attention. Some studies have focused on Arabic AES, utilizing feature engineering and applying machine learning to develop models Alghamdi et al. (2014); Al-Shalabi (2016); Alobed et al. (2021); Gaheen et al. (2021). This work and related research has only partially addressed the key challenges hindering the development of Arabic AES models, namely, the lack of publicly available large-scale annotated datasets designed to improve the quality of Arabic writing. Ghazawi and Simpson (2024) introduced AR-AES, a benchmark dataset that contains 2,046 essays written by undergraduate students across three different university faculties, featuring a mix of question types. Two educators from each faculty annotated and graded the essays using clearly defined rubrics, aiming to assess students’ academic performance across various disciplines. Unlike AR-AES, our work aims to evaluate and assess writing proficiency using the CEFR rubrics as these constitute an internationally accepted standard. QAES Bashendy et al. (2024) presented the first publicly available trait-specific annotations for Arabic AES. QAES is based on the work of the Qatari Corpus of Argumentative Writing (QCAW) Ahmed et al. (2024), which consists of 195 Arabic argumentative essays. They proposed a multi-layer of annotation traints such as coherence, organization, grammar, and others. Despite its excellent annotation, it is small in size and limited to two prompts. While QAES multi-traits scores are publicly available, the QCAW holistic score is not public. Habash and Palfreyman (2022) presented the Zayed University Arabic-English Bilingual Undergraduate Corpus (ZAEBUC). This corpus comprises non-parallel essays in Arabic and English related to three prompts collected from first-year university students with differing writing proficiency. ZAEBUC includes 216 annotated Arabic essays featuring manual annotations for syntactic and morphological characteristics and a CEFR-based proficiency assessment. Again, ZAEBUC is small in size and limited to three prompts. Researchers have investigated several approaches to data augmentation such as sampling techniques, introducing noise into texts, or paraphrasing, to overcome the challenges of data scarcity and quality Li et al. (2022). The recent development of LLMs has paved the way for researchers to explore promising new data synthesis solutions Wang et al. (2024a); Long et al. (2024). Transformers and LLMs can closely mirror real-world distributions while introducing valuable variations across multiple tasks and domains Wang et al. (2024b). GPT models such as ChatGPT, GPT-4, and GPT-4o demonstrate significant capabilities in augmenting and generating synthetic essays for the purpose of English AES Ramesh and Sanampudi (2022). LLMs and Transformer models with Arabic capabilities have shown robust generalization across different Arabic NLP tasks such as: Question-Answering Samuel et al. (2024), Code-Switching Alharbi et al. (2024), Name Entity Recognition (NER) Sabty et al. (2021), Grammar Error Correction (GEC) Solyman et al. (2023), and Sentiment Analysis Refai et al. (2023). However, to the best of our knowledge no research has utilized such models to generate Arabic essays across CEFR writing proficiency levels.
\<وسائل التواصل الاجتماعي وتأثيرها على الفرد والمجتمع.>
How do social media affect individuals and society?
\<كيف نعزز ثقافة التسامح في المجتمع؟>
How can the UAE promote a culture of tolerance in society?
\<التطور الحضاري الذي تشهده دولة الإمارات العربية المتحدة>
What do you think are the most important developments in the UAE at the moment?
Table 1: The prompts given to the essay writers in the ZAEBUC corpus Habash and Palfreyman (2022).
3 Data: The ZAEBUC Corpus
For all our experiments and reported results, we use the ZAEBUC corpus Habash and Palfreyman (2022). ZAEBUC comprises essays written by native Arabic speakers, which were manually corrected and annotated for writing proficiency using the CEFR Council of Europe (2001) rubrics and scale. Each essay was independently annotated by three CEFR-proficient bilingual speakers. Habash and Palfreyman (2022), assigned a holistic CEFR level to each essay by converting the three CEFR ratings into numerical scores (ranging from 1 to 6) and then taking the rounded average. The essays in the corpus were limited to three prompt choices on Social Media, Tolerance, and Development; see Table 1. We use the splits created by Alhafni et al. (2023). Table 2 presents the ZAEBUC corpus CEFR level distributions in terms of the holistic CEFR levels.
CEFR Level Count Percentage
A1 0 0%
A2 7 3%
B1 110 51%
B2 80 37%
C1 11 5%
C2 0 0%
Unassessable 6 3%
Total 214 100%
Table 2: ZAEBUC corpus CEFR level distributions.
4 Synthetic Data Augmentation
We propose a synthetic data augmentation approach leveraging the ZAEBUC dataset to generate synthetic essays that align with CEFR rubrics and have features similar to human text. The pipeline utilizes three phases: Building Essay Prompts, Feature Profiling, and finally Data Augmentation.
4.1 Building Essay Prompts
We began by compiling a collection of essay prompts covering diverse categories and proficiency levels. We considered three proficiency levels aligned with the CEFR scale: Beginner (A1–A2), Intermediate (B1–B2), and Advanced (C1–C2). While some categories, such as Hobbies, are suitable for all levels, others, like Politics, Technology, and Education, were introduced specifically for advanced learners. Using various LLMs, including GPT-4o, Gemini, and Copilot, we generated 100 prompts across these categories, ensuring relevance for Arabic essay writing and an unbiased distribution across proficiency levels. We then manually reviewed the prompts, removing redundancies and irrelevant entries while maintaining comprehensive coverage. The final collection consists of 152 balanced and diverse prompts. Table 3 presents the selected categories and the distribution of prompts across levels, while Table 4 provides example prompts from the Hobbies category.
Topic B I A
Culture and Traditions 1 3 2
Daily Life 2 2 2
Education 3 6 8
Environment 2 2 3
Future 1 2 2
History and Culture 2 2 2
Hobbies 3 2 2
Imaginary 5 2 2
Life/Time Management 4 4 2
Personal Experiences 7 2 2
Relations 4 2 2
School Life 4 2 2
Sport and Health 2 3 1
Technology and Media 2 8 6
Travel and Experience 1 2 1
Politics and Government 2 2 7
Social Issues 2 7 6
Total 47 53 52
Table 3: Count of Arabic text prompts by level and topic. B: Beginner level (A1, A2), I: Intermediate level (B1, B2), A: Advanced level (C1, C2).
Level Arabic Prompt English Prompt
Beginner \<ما هي هوايتك المفضلة؟ >  $\bullet$ $\bullet$ What is your favorite hobby?
\<تحدث عن نشاط تحبه في عطلة نهاية الأسبوع.>  $\bullet$ $\bullet$ Talk about a weekend activity you love.
\< ما هي الرياضة التي تحب ممارستها؟ لماذا؟>  $\bullet$ $\bullet$ What is your favorite sport? Why?
Intermediate \<في ممارستها؟ لماذا تحبها؟> \<ما هي هوايتك المفضلة وكيف بدأت >   $\bullet$ $\bullet$ What is your favorite hobby and how did you start practicing it? Why do you enjoy it?
\< ما هي ولماذا تهمك؟> \< هل ترغب في تعلم هواية جديدة؟>   $\bullet$ $\bullet$ Do you wish to learn a new hobby? What is it and why does it interest you?
Advanced \<ناقش كيف تؤثر الهواية على صحتك النفسية والجسدية.>   $\bullet$ $\bullet$ Discuss how hobbies impact your mental and physical health.
\<كيف يمكن أن تكون الهوايات وسيلة للتعبير عن الذات؟> How can hobbies serve as a means of self-expression?
\< هل هناك هواية جديدة ترغب في تجربتها؟ >   $\bullet$ $\bullet$ Is there a new hobby you wish to try? Discuss the reasons
\<ناقش الأسباب التي تجعلك مهتمًا بها وكيف تعتقد أنها ستفيدك.> you are interested in it and how you believe it will benefit you.
Table 4: Examples of prompts related to the topic of Hobbies and classified into one of three different levels.
4.2 Feature Profiling
We construct linguistic profiles for each CEFR level using the ZAEBUC corpus. Each profile contains various levels of linguistic information. Representing different lexical and syntactic features, we use the number of words/sentences ($N_{w}$,$N_{s}$), the number of tokens /vocabulary ($N_{v}$), words/sentences lengths ($L_{w}$,$L_{s}$), and sentence complexity measured by syntactic tree depth ($D_{s}$). We define the lexical diversity (Type-Token Ratio, TTR) as:
TTR=Unique TokensTotal $\text{TTR}=\frac{\text{Unique Tokens}}{\text{Total Tokens}}$ (1)
Similarly, we calculate the sentence complexity by:
$C_{s}=\frac{\sum_{i=1}^{N}D_{i}}{N}$ (2)
where $D_{i}$ is the syntactic depth of sentence $i$ and $N$ is the total number of sentences. For morphological features, we use the ZAEBUC morphological annotations: the most frequent POS tags, such as nouns, verbs, adjectives, etc. We aggregate all extracted features across the essays to get a quantitative representation at different writing CEFR levels, which serves as a reference for later stages.
4.3 Zero-shot Data Augmentation
Effective LLM prompt engineering has become increasingly important, as the model’s output varies based on the prompt, provided instructions, and prompt language. Previous studies in Arabic NLP have shown that using English as the instruction language for input prompts can improve output quality Kmainasi et al. (2024); Koto et al. (2024). In our approach, we experiment with various prompts for zero-shot data augmentation to identify those that produce human-like text while adhering to guideline instructions. We use GPT-4o as our generation model due to its affordability and larger token capacity for both input and output. The GPT prompts include: (a) the target CEFR level, (b) CEFR guidelines and instructions, (c) the linguistic profile for the targeted CEFR level to control the prompt output, and (d) the topic prompt or question from the previously mentioned topic prompts dataset. To check the quality of the generated essays and whether they follow the prompt instructions, we build a linguistic feature profile for each augmented essay. We employ the cosine similarity to assess the faithfulness of the augmented approach and whether the augmented essays are close to the original CEFR profiles. We applied the cosine similarity as in equation 3 either to one of the features profiles (lexcial/syntactical features , morphological features or both of them).
$\cos(\theta)=\frac{\sum_{i}P_{i}Q_{i}}{\sqrt{\sum_{i}P_{i}^{2}}\cdot\sqrt{\sum_{i}Q_{i}^{2}}}$ (3)
where $P_{i}$ and $Q_{i}$ are the probability values of linguistic features in the original and generated distribution profiles. This metric ensures that the synthetic data closely aligns with real human essay patterns. Later we calculate the matching agreements by:
$\text{Agreement}=\frac{\sum_{i=1}^{n}(\hat{y}_{i}=y_{i})}{n}$ (4)
where $\hat{y}_{i}$ is the predicted level and $y_{i}$ is the ground truth. We conducted multiple rounds of prompt engineering refinements to improve the quality of the generated Arabic essays and ensure alignment with CEFR levels. First, we found that straightforward prompts without explicit controlled linguistic instructions and explanations resulted in incoherent essays, including irrelevant topics and English text, achieving only 20.5% matching agreement with linguistic feature profiles. In a subsequent round, we introduced detailed definitions of linguistic features and restricted outputs to Arabic-only text, which improved agreement to 26%. However, the model still occasionally produced incomplete essays and injected text from the prompt into the essay. The most effective prompt structure format is illustrated in Figure 1. We separated system-level control instructions from user-defined parameters, thereby providing clearer guidance for structured and proficiency-aligned text generation. This refinement increased agreement to 27.5%, demonstrating that precisely controlled instructions enhance LLM performance in structured writing tasks. Ultimately, we generated 3,040 Arabic essays covering all CEFR levels and various topics, where each prompt produced ten essays. The structured and controlled prompt design not only improved alignment with learner writing styles but also established a consistent framework for synthetic Arabic text generation. Table 5 presents the distribution of generated essays across different CEFR levels.
CEFR Level Count Percentage
A1 470 15.5%
A2 470 15.5%
B1 530 17.4%
B2 530 17.4%
C1 520 17.1%
C2 520 17.1%
Total 3,040 100%
Table 5: The generated corpus CEFR level distributions.
Figure 1: GPT-4o prompts messages that has been used to generate Arabic essays
5 Error Injection
Human-generated text naturally contains some grammatical errors and linguistic infelicities. In order to generate human-like essays, we need to include similar kinds of errors that reflect the level of writing attainment. In this phase, we prompt GPT-4o to inject errors into the previously generated essays while maintaining their aligned CEFR levels by utilizing error profiling.
5.1 Error Profiling
Error Distribution Profiles
To model the distribution of errors to inject into the synthetic essays, we again leverage the ZAEBUC corpus, which contains both the erroneous essays aligned with the manually corrected ones. We followed the same methodology we used to construct the linguistic feature profiles for each CEFR level to develop error distribution profiles aligned with CEFR levels. The error profile captures and reflects the authentic distribution patterns observed in human writing at different CEFR levels. Developing an Error Instruction Repository To prompt GPT-4o to generate essays containing errors we applied the GED model proposed by Alhafni et al. (2023) to the ZAEBUC corpus to annotate errors using 13 error tags and to obtain error distributions for each CEFR level. We created the repository using the same 13 tags (ReplaceM, ReplaceO, ..), where we also added a formal definition of what those tags describe in terms of linguistic errors. In addition, we expanded the error taxonomy by splitting it into finer-grained classes. Each error instruction is followed by an example showing the correct word and the erroneous version. The explanation was based on the extended ALC taxonomy Alfaifi et al. (2013), which was refined later and introduced as ARETA Belkebir and Habash (2021). Figure 2 shows some examples from our error instruction repository.
Figure 2: an Example of orthographical error instructions from the developed errors instructions repository
5.2 GPT-Based Error Injection
We prompted GPT-4o to inject errors into the synthetically generated essays based on the error distribution profiles while maintaining the CEFR level. The model processed one essay at a time in a zero-shot setting, except that we included the definition and explanation of the error tags. For example, Replace_M indicates a morphological error, while Merge targets two mistakenly split tokens that need to be merged, and so on. After conducting multiple experiments, we observed the following issues: (i) The model struggled to follow the predefined error distribution perhaps due to the complexity of the prompts. (ii) The model was confused by certain error tags, particularly SPLIT, MERGE as well as the UNK (unknown errors). These errors were mainly ignored in the injected text. (iii) We calculated the cosine similarity between the main error profile and the injected essays’ error distributions as shown in equation 3. When we injected all errors at once, the similarity agreements did not exceed 20%; however, when we reduced the number of error tags per essay or omitted the misleading tags, the agreements significantly improved, reaching 86% Therefore we implemented a method where each error type was injected separately. This required multiple iterations over the same essay, corresponding to the number of error tags shown in the error distribution profile for each CEFR level. Figure 3 shows an example of a GPT-4o prompt for error injection. Some errors occurred more frequently than others among Arabic writers. For example, orthographic errors are more frequent than semantic errors. So, rather than repeating a single error instruction for all instances of an error tag, we randomly select a weighted number of error instructions proportional to the average occurrence of each error type. Figure 4 shows the pseudocode for the selection process. The full pseudocode is in Appendix A.
Figure 3: Sample GPT-4o error injection prompt
Figure 4: Error injection based on average error count
5.3 Controlled Error Injection
We introduce a controlled method for injecting errors into clean text, ensuring that the resulting erroneous sentences follow the empirical error distributions observed at each CEFR level. More formally, given an input sentence ($X$) and its CEFR level ($L$), we introduce errors in two steps: Error Type Prediction and Error Realization.
Error Type Prediction
We estimate the probability of an error type occurring at a given word, i.e., $P(error\_type|word)$. To do so, we leverage ARETA in a reverse annotation process where we process correct–erroneous sentence pairs, tagging each correct word with its corresponding error type. Using this annotated data, we train a token-level BERT classifier to predict the most likely error type for each word in a given correct sentence. We fine-tune CAMeLBERT MSA Inoue et al. (2021) to build our classifier.
Error Realization
To determine how a word should be corrupted, we first align correct–erroneous sentence pairs using the algorithm proposed by Alhafni et al. (2023). For each aligned pair, we extract edit transformations that capture the operations required to convert a correct word into its erroneous counterpart. Using this data, we estimate $P(transformation|error\_type)$ with a bigram Maximum Likelihood Estimation (MLE) lookup model: $count(transformation,error\_type)$ / $count(error\_type)$. During inference, we apply the BERT classifier to predict error types for each word in a sentence. We then filter these predictions, retaining only error types relevant to the sentence’s CEFR level. Finally, the MLE model selects the most probable corruption for a given error type. A complete example of a B1–level essay generated by the proposed model is in Figure 5.
Figure 5: An Example of a B1 Arabic Essay generated by GPT-4o using the Hobbies prompt and the same essay after injecting errors by the controlled BERT-based model.
Average Reference Multi-Reference
Train Data QWK Acc F1 P R QWK Acc F1 P R
Zaebuc-1 22.44 57.58 24.50 23.33 26.76 61.06 84.85 43.70 42.50 45.31
Zaebuc-1 + GPT essays 14.92 60.61 26.43 25.55 27.45 96.00 96.97 92.32 98.04 88.89
Zaebuc-1 + BERT errors 27.87 57.58 38.02 35.86 44.93 82.70 87.88 71.66 70.83 74.38
Zaebuc-1 + GPT errors_1 17.14 57.58 25.64 25.18 26.27 96.47 96.97 94.16 97.92 91.67
Zaebuc-1 + GPT errors_0 20.84 57.58 32.76 31.53 46.08 93.79 93.94 95.12 96.49 94.44
Table 6: Performance comparison of different training datasets. GPT essays are the original correct essays generated from GPT-4o, BERT errors are the erroneous essays using the controlled injection BERT model, GPT errors_1 are the erroneous essays using GPT-4o with one shot error example, while GPT errors_0 with Zero shot settings
6 Evaluation
6.1 Metrics
Our primary evaluation metric is Quadratic Weighted Kappa (QWK) Cohen (1968), the most widely-used metric in recent AES research Ke and Ng (2019). We also report accuracy, macro precision (P), recall (R), and F1 scores. Model predictions are evaluated against the average of the gold holistic CEFR labels and in a multi-reference setting against the three individual gold labels per essay (§3).
6.2 Data
We use the ZAEBUC dataset for all the experiments. The dataset was split (70%,15%,15%) for Train, Dev and Test sets, respectively Alhafni et al. (2023). Table 6 shows the performance results of the BERT-based Scoring Model utilizing different training sets.
6.3 Model
We treat the task of automated essay scoring (AES) as a text classification problem. For Arabic AES, we fine-tune CAMeLBERT MSA Inoue et al. (2021) on the training split of ZAEBUC. The models were trained by using the average CEFR gold labels. During training, we ignore the essays that are labeled as Unassessable, but we penalize the models for missing them in the evaluation. We fine-tune the models for 5 epochs, with a maximum sequence length of 512, a learning rate of 5e-5, and a batch size of 32.
6.4 Results
Our results are presented in Table 6. Our baseline system, only trained on the ZAEBUC training set, indicates room for improvement, with the F1 at 24.50% and QWK at 22.44%. We then switched between different datasets to enhance the Arabic AES model and to measure the impact of data augmentation on the model. We introduced the 3,040 correct essays generated using GPT-4o, which dropped the QWK compared to the baseline while increasing the F1. However, we observed significant improvements in the multi-reference settings for the QWK (96.00%) and F1 (92.32%). Since this is an error-free, synthetic dataset, we further refine the model by introducing essays that mimic human-like writing styles. We compare the two approaches introduced in section 5: (1) GPT-based error injection (with and without error instruction examples) and (2) the controlled error injection method based on CAMeLBERT MSA. Inoue et al. (2021). The results demonstrate that the controlled error model enhances performance across all metrics, particularly in the average-reference setting, that achieved 27.87 % and 38.02 for QWK and F1 respectively. However, the GPT-based error injection performs best under the multi-reference setting, where the one-shot essays increased the QWK to 96.47 and the zero-shot erroneous essays improved the F1 to 95.12%. These results suggest that LLMs and Transformer methods generalize effectively in AES models and also improve performance. The results emphasize that a controlled methodology is more reliable in mimicking patterns of learner errors and mistakes. This positions the controlled BERT-based model as a vital component of the data augmentation pipeline, effectively addressing the limitations of the GPT models.
7 Conclusions and Future Work
This paper proposes a hybrid framework for building an Arabic AES system. We utilize both LLMs and Transformers to overcome the challenges of dataset scarcity by generating synthetic data that mimics the writing styles of Arabic learners. We leverage the ZAEBUC corpus and build linguistic feature profiles as well as error profiles that align with each CEFR score. We prompt GPT-4o to generate 3,040 Arabic essays based on multiple CEFR levels corresponding to 152 prompts. However, GPT-4o’s effectiveness depends heavily on prompt engineering, achieving only 27.5% agreement with the original linguistic feature profiles. To introduce errors into GPT-generated essays, we compare two approaches: (1) GPT-4o error injection, which struggles with realistic learner writing due to misinterpreted linguistic errors or complex prompts, causing inconsistencies; (2) a controlled method fine-tuning the CAMeLBERT MSA model to inject errors proportionally to their natural occurrence. The latter method is an essential component of data augmentation for Arabic AES. Our evaluation, using a fine-tuned BERT classifier, shows that the proposed hybrid framework that combines GPT-generated data and a controlled error injection method shows consistent improvement over the baseline (QWK: 27.87%, F1: 38.02%), achieving more predictable and interpretable results. These results suggest the controlled approach to generating learner errors better captures the actual distribution of errors across proficiency levels. In future work, we plan to expand the dataset by incorporating real essay samples, which will enhance the linguistic and error profiles for each CEFR level, thereby improving the proposed data augmentation model. In addition, we will deploy the AES system as part of an interactive educational writing assistant tool to help students enhance their Arabic writing with immediate feedback on errors and CEFR scores.
8 Limitations
Despite the effectiveness of our hybrid Arabic AES framework, we note several limitations related to the quality of generated Arabic essays, error injection accuracy, and the generalization of the AES model. The lack of A1 and C2 essays in ZAEBUC means that there is no gold reference data for these levels, which may impact both linguistic and error profiles, affecting the accuracy of GPT-generated essays. Furthermore, different biases exist in both ZAEBUC and the GPT outputs: ZAEBUC overrepresents B-level essays, skewing model training; and GPT-4o tends to favor Arabic culture themes, thereby limiting diversity in the generated texts, and usability in other contexts. In addition, due to the lack of comprehensive gold data, GPT struggles to fully replicate real learner writing styles, achieving only 27.5% agreement with linguistic feature profiles. Another limitation is the model’s ability to generalize across various domains and question types. The AES system may struggle with broader writing tasks and alternative prompts since the dataset and augmentation methods focus on predefined prompts. Relying solely on CEFR as a holistic scoring method limits interpretability. Enhancing the dataset with multi-trait annotations, such as coherence, argumentation, and organization, could improve scoring accuracy and feedback quality. Moreover, better-controlled GPT prompting could refine the quality and diversity of generated essays, reducing biases and improving alignment with real learner writing patterns.
9 Ethical Considerations
While Arabic AES systems provide significant support in assessing Arabic learners’ writing proficiency, it is essential to highlight the ethical implications of their use. Automatic assessment and scoring may lead to misjudgments that could distress learners and students, especially if their work is incorrectly evaluated. AES tools should serve as an educational assistive technology, complementing the teacher’s judgment, not replacing it in educational settings.
Appendix A GPT-4o Error injection algorithm
Figure 6: GPT-4o Error injection and verification algorithm

### v2

@@
 (2021).
-Unlike for English, for Modern Standard Arabic (MSA) it is difficult to develop robust and scalable AES systems, primarily due to the lack of essay datasets necessary for building effective Arabic AES Lim et al.
+Unlike for English, it is difficult to develop robust and scalable AES systems for Modern Standard Arabic (MSA), primarily due to the lack of essay datasets necessary for building effective Arabic AES Lim et al.
 (2021); Elhaddadi et al.
 (2024).
-This paper presents our proposed framework to tackle the issue of data scarcity and quality by utilizing Transformers and Large Language Models (LLMs) to generate and build a synthetic dataset.
+This paper presents a framework to tackle the issue of data scarcity and quality by utilizing Transformers and Large Language Models (LLMs) to generate and build a synthetic dataset.
 Our approach begins with prompting GPT-4o to generate a variety of Arabic essays covering multiple topics and different writing proficiency levels as defined by the Common European Framework of Reference (CEFR) Council of Europe (2001).
@@
   • Proposing a framework based on LLMs and Transformers for augmenting Arabic essays that accurately reflect human writing patterns.
-  • Creating a synthetic Arabic AES dataset with 3,040 essays annotated with CEFR proficiency levels. [^1] The full dataset will be publicly available for research.
-  • Developing an Arabic essay auto-marking system using a BERT-based scoring model, enabling accurate and scalable evaluation of Arabic essays based on CEFR standards.
-The rest of the paper is organized as follows: §2 reviews related work on AES, §3 describes the dataset, and §4 outlines our data augmentation approach.
-§5 details the error injection methods, followed by evaluation in §6 and §7 presents conclusion and future work.
+  • Creating a synthetic Arabic AES dataset with 3,040 essays annotated with CEFR proficiency levels.
+  • Developing an Arabic AES system using a BERT-based model, enabling accurate and scalable evaluation of Arabic essays based on CEFR standards.
+The rest of the paper is organised as follows: §2 reviews related work on AES, §3 describes the dataset, and §4 outlines our data augmentation approach.
+§5 details the error injection methods, followed by an evaluation in §6.
+We discuss our results in §7 and §8 presents the conclusion and future work.
 2 Related Work
-The field of AES has been investigated extensively, particularly in English Lim et al.
-(2021); Ramesh and Sanampudi (2022), where some valuable tools have been introduced such as IntelliMetric Elliott et al.
+AES has been investigated extensively, particularly in English Lim et al.
+(2021); Ramesh and Sanampudi (2022), where multiple tools have been introduced such as IntelliMetric Elliott et al.
 (2003), e-rater Attali and Burstein (2006), Grammarly, [^2] https://app.grammarly.com/ Write and Improve [^3] https://writeandimprove.com/ Yannakoudakis et al.
 (2018), and others.
-The development and validation of English AES have been enabled by the large-scale annotated datasets that have been released, such as the First Cambridge English (FCE) dataset [^4] https://ilexir.co.uk/datasets/index.html Yannakoudakis et al.
+The development of English AES systems has been enabled by large scale annotated datasets such as the First Cambridge English (FCE) dataset [^4] https://ilexir.co.uk/datasets/index.html Yannakoudakis et al.
 (2011), Automated Student Assessment Prize (ASAP) dataset, [^5] https://www.kaggle.com/c/asap-aes the TOEFL11 corpus Blanchard (2013), and the ICLE (International Corpus of Learner English) Granger (2003).
-These datasets contained thousands of student essays with proficiency-level grades, often along multiple dimensions.
-In contrast, research on Arabic AES has not received comparable attention.
-Some studies have focused on Arabic AES, utilizing feature engineering and applying machine learning to develop models Alghamdi et al.
+These datasets contain thousands of student essays with proficiency level grades, often along multiple dimensions.
+In contrast, Arabic AES research has received less attention.
+Some studies have applied feature engineering and machine learning to develop models Alghamdi et al.
 (2014); Al-Shalabi (2016); Alobed et al.
 (2021); Gaheen et al.
-(2021).
-This work and related research has only partially addressed the key challenges hindering the development of Arabic AES models, namely, the lack of publicly available large-scale annotated datasets designed to improve the quality of Arabic writing.
-Ghazawi and Simpson (2024) introduced AR-AES, a benchmark dataset that contains 2,046 essays written by undergraduate students across three different university faculties, featuring a mix of question types.
-Two educators from each faculty annotated and graded the essays using clearly defined rubrics, aiming to assess students’ academic performance across various disciplines.
-Unlike AR-AES, our work aims to evaluate and assess writing proficiency using the CEFR rubrics as these constitute an internationally accepted standard.
-QAES Bashendy et al.
-(2024) presented the first publicly available trait-specific annotations for Arabic AES.
-QAES is based on the work of the Qatari Corpus of Argumentative Writing (QCAW) Ahmed et al.
+(2021), but they partially address key challenges, especially the scarcity of large, publicly available annotated datasets for improving Arabic writing quality.
+Ghazawi and Simpson (2024) introduced AR-AES, a benchmark of 2,046 undergraduate essays from three university faculties, annotated by two educators per faculty using rubrics to assess academic performance.
+In contrast, our work focuses on writing proficiency, using the CEFR standard.
+Bashendy et al.
+(2024) presented QAES, the first publicly available trait specific annotations for Arabic AES.
+QAES extends the Qatari Corpus of Argumentative Writing (QCAW) Ahmed et al.
 (2024), which consists of 195 Arabic argumentative essays.
-They proposed a multi-layer of annotation traints such as coherence, organization, grammar, and others.
-Despite its excellent annotation, it is small in size and limited to two prompts.
-While QAES multi-traits scores are publicly available, the QCAW holistic score is not public.
+They implemented multi-layered annotation of traits such as coherence, organization, grammar, and others.
+Despite its comprehensive annotation, it is small in size and limited to two prompts.
+While QAES multi-traits scores are publicly available, the QCAW holistic score is not.
 Habash and Palfreyman (2022) presented the Zayed University Arabic-English Bilingual Undergraduate Corpus (ZAEBUC).
@@
 Again, ZAEBUC is small in size and limited to three prompts.
-Researchers have investigated several approaches to data augmentation such as sampling techniques, introducing noise into texts, or paraphrasing, to overcome the challenges of data scarcity and quality Li et al.
+Researchers have explored data augmentation methods like sampling, noise injection, and paraphrasing to address data scarcity and quality Li et al.
 (2022).
@@
 (2024b).
-GPT models such as ChatGPT, GPT-4, and GPT-4o demonstrate significant capabilities in augmenting and generating synthetic essays for the purpose of English AES Ramesh and Sanampudi (2022).
-LLMs and Transformer models with Arabic capabilities have shown robust generalization across different Arabic NLP tasks such as: Question-Answering Samuel et al.
-(2024), Code-Switching Alharbi et al.
-(2024), Name Entity Recognition (NER) Sabty et al.
-(2021), Grammar Error Correction (GEC) Solyman et al.
+GPT models have shown strong capabilities in generating synthetic essays for English AES Ramesh and Sanampudi (2022).
+LLMs and Transformers have also generalized well in Arabic NLP tasks, including Question Answering Samuel et al.
+(2024), Code Switching Alharbi et al.
+(2024), NER Sabty et al.
+(2021), Grammatical Error Correction Alhafni and Habash (2025); Solyman et al.
 (2023), and Sentiment Analysis Refai et al.
@@
 However, to the best of our knowledge no research has utilized such models to generate Arabic essays across CEFR writing proficiency levels.
+3 Data: The ZAEBUC Corpus
+For all our experiments, we use the ZAEBUC corpus Habash and Palfreyman (2022).
+ZAEBUC comprises essays written by native Arabic speakers, which were manually corrected and annotated for writing proficiency using the CEFR Council of Europe (2001) rubrics and scale.
+Each essay was annotated by three CEFR-proficient bilingual speakers.
+Habash and Palfreyman (2022), assigned a holistic CEFR level to each essay by converting the three CEFR ratings into numerical scores (ranging from 1 to 6) and then taking the rounded average.
+The essays in the corpus were limited to three prompt choices on Social Media, Tolerance, and Development; see Table 1.
+We use the splits created by Alhafni et al.
+(2023).
+Table 2 shows the CEFR level distribution of the ZAEBUC corpus based on holistic CEFR scores.
+The ZAEBUC corpus is limited in size and skewed toward B1–B2 levels, with no A1 or C2 essays.
+This common imbalance in Arabic learner data motivated our synthetic approach to create a more balanced CEFR distribution.
 \<وسائل التواصل الاجتماعي وتأثيرها على الفرد والمجتمع.>
@@
 Table 1: The prompts given to the essay writers in the ZAEBUC corpus Habash and Palfreyman (2022).
-3 Data: The ZAEBUC Corpus
-For all our experiments and reported results, we use the ZAEBUC corpus Habash and Palfreyman (2022).
-ZAEBUC comprises essays written by native Arabic speakers, which were manually corrected and annotated for writing proficiency using the CEFR Council of Europe (2001) rubrics and scale.
-Each essay was independently annotated by three CEFR-proficient bilingual speakers.
-Habash and Palfreyman (2022), assigned a holistic CEFR level to each essay by converting the three CEFR ratings into numerical scores (ranging from 1 to 6) and then taking the rounded average.
-The essays in the corpus were limited to three prompt choices on Social Media, Tolerance, and Development; see Table 1.
-We use the splits created by Alhafni et al.
-(2023).
-Table 2 presents the ZAEBUC corpus CEFR level distributions in terms of the holistic CEFR levels.
 CEFR Level Count Percentage
@@
 4.1 Building Essay Prompts
-We began by compiling a collection of essay prompts covering diverse categories and proficiency levels.
-We considered three proficiency levels aligned with the CEFR scale: Beginner (A1–A2), Intermediate (B1–B2), and Advanced (C1–C2).
-While some categories, such as Hobbies, are suitable for all levels, others, like Politics, Technology, and Education, were introduced specifically for advanced learners.
-Using various LLMs, including GPT-4o, Gemini, and Copilot, we generated 100 prompts across these categories, ensuring relevance for Arabic essay writing and an unbiased distribution across proficiency levels.
-We then manually reviewed the prompts, removing redundancies and irrelevant entries while maintaining comprehensive coverage.
+We began by compiling a diverse set of essay prompts across various categories and CEFR levels.
+While not directly drawn from established frameworks, our prompts were inspired by themes common in language assessments, including placement tests and academic writing.
+We aimed to cover familiar and level-appropriate topics, such as social issues, education, and personal experiences, while ensuring balance across the CEFR bands.
+We considered three proficiency levels: Beginner (A1–A2), Intermediate (B1–B2), and Advanced (C1–C2).
+General themes, such as hobbies, suited all levels, while more complex topics, including politics, Technology, and Education, were reserved for advanced learners.
+Using LLMs like GPT-4o [^6] https://openai.com/index/hello-gpt-4o/, Gemini [^7] https://gemini.google.com/app, and Copilot [^8] https://copilot.microsoft.com/, we generated 100 prompts, followed by a manual review to remove redundancies and ensure both relevance for Arabic essay writing and balanced proficiency coverage.
 The final collection consists of 152 balanced and diverse prompts.
-Table 3 presents the selected categories and the distribution of prompts across levels, while Table 4 provides example prompts from the Hobbies category.
+Table 3 presents the selected categories and the distribution of the prompts across levels, while Table 4 provides example prompts for the Hobbies category.
 Topic B I A
@@
 We use GPT-4o as our generation model due to its affordability and larger token capacity for both input and output.
-The GPT prompts include: (a) the target CEFR level, (b) CEFR guidelines and instructions, (c) the linguistic profile for the targeted CEFR level to control the prompt output, and (d) the topic prompt or question from the previously mentioned topic prompts dataset.
-To check the quality of the generated essays and whether they follow the prompt instructions, we build a linguistic feature profile for each augmented essay.
-We employ the cosine similarity to assess the faithfulness of the augmented approach and whether the augmented essays are close to the original CEFR profiles.
-We applied the cosine similarity as in equation 3 either to one of the features profiles (lexcial/syntactical features , morphological features or both of them).
+The GPT prompts include (a) the target CEFR level, (b) CEFR guidelines and instructions, (c) the linguistic profile for the targeted CEFR level to control the prompt output, and (d) the topic prompt or question from the previously mentioned topic prompts dataset.
+For these missing levels (A1 and C2), instead of injecting a pre-defined profile, GPT-4o was directly prompted to act as an assistant and generate data based on the general standards and rubrics of the CEFR.
+To check the quality of the generated essays and whether they follow the prompt instructions, we build a linguistic feature profile (vector) for each augmented essay.
+We then assess the alignment between the generated essays and the reference CEFR-level profiles by computing their feature vectors’ cosine similarity as in equation 3.
+Specifically, given two real-valued feature vectors $P_{i}$ (the CEFR reference profile) and $Q_{i}$(the generated essay), the cosine similarity is calculated as:
 $\cos(\theta)=\frac{\sum_{i}P_{i}Q_{i}}{\sqrt{\sum_{i}P_{i}^{2}}\cdot\sqrt{\sum_{i}Q_{i}^{2}}}$ (3)
-where $P_{i}$ and $Q_{i}$ are the probability values of linguistic features in the original and generated distribution profiles.
 This metric ensures that the synthetic data closely aligns with real human essay patterns.
-Later we calculate the matching agreements by:
+Based on the computed similarity score, we assign a predicted CEFR level to each essay.
+Later, we calculate the alignment between the predicted CEFR level and the target level specified in the GPT-4o prompt (ground truth) using the following agreement formula:
 $\text{Agreement}=\frac{\sum_{i=1}^{n}(\hat{y}_{i}=y_{i})}{n}$ (4)
 where $\hat{y}_{i}$ is the predicted level and $y_{i}$ is the ground truth.
+This process evaluates how well GPT-4o succeeded in aligning the generated content with the intended proficiency level, serving as a measure of agreement rather than a prediction from an external model.
 We conducted multiple rounds of prompt engineering refinements to improve the quality of the generated Arabic essays and ensure alignment with CEFR levels.
@@
 This refinement increased agreement to 27.5%, demonstrating that precisely controlled instructions enhance LLM performance in structured writing tasks.
-Ultimately, we generated 3,040 Arabic essays covering all CEFR levels and various topics, where each prompt produced ten essays.
-The structured and controlled prompt design not only improved alignment with learner writing styles but also established a consistent framework for synthetic Arabic text generation.
+Ultimately, we generated 3,040 Arabic essays covering all CEFR levels and various topics, where each prompt was used to create ten essays.
+This effort was intentionally designed to address the imbalanced CEFR distribution in the original ZAEBUC corpus, where B-level essays were overrepresented.
+By constructing a more balanced synthetic dataset, we aimed to enhance model performance across the full proficiency spectrum.
+The structured and controlled prompt design also improved alignment with learner writing styles while providing a consistent framework for generating realistic Arabic essays.
 Table 5 presents the distribution of generated essays across different CEFR levels.
+The full dataset statistics are provided in Appendix A.1.
 CEFR Level Count Percentage
@@
 Table 5: The generated corpus CEFR level distributions.
-Figure 1: GPT-4o prompts messages that has been used to generate Arabic essays
+Figure 1: GPT-4o prompts messages that have been used to generate Arabic essays
 5 Error Injection
 Human-generated text naturally contains some grammatical errors and linguistic infelicities.
-In order to generate human-like essays, we need to include similar kinds of errors that reflect the level of writing attainment.
+In order to create human-like essays, we need to add similar kinds of errors to the synthetic essays that reflect the level of writing attainment.
 In this phase, we prompt GPT-4o to inject errors into the previously generated essays while maintaining their aligned CEFR levels by utilizing error profiling.
@@
 Error Distribution Profiles
-To model the distribution of errors to inject into the synthetic essays, we again leverage the ZAEBUC corpus, which contains both the erroneous essays aligned with the manually corrected ones.
+To model the distribution of errors to inject into the synthetic essays, we again leverage the ZAEBUC corpus, which contains the erroneous essays aligned with the manually corrected ones.
 We followed the same methodology we used to construct the linguistic feature profiles for each CEFR level to develop error distribution profiles aligned with CEFR levels.
 The error profile captures and reflects the authentic distribution patterns observed in human writing at different CEFR levels.
-Developing an Error Instruction Repository To prompt GPT-4o to generate essays containing errors we applied the GED model proposed by Alhafni et al.
+Developing an Error Instruction Repository To prompt GPT-4o to generate essays containing errors we applied the Grammatical Error Detection (GED) model proposed by Alhafni et al.
 (2023) to the ZAEBUC corpus to annotate errors using 13 error tags and to obtain error distributions for each CEFR level.
-We created the repository using the same 13 tags (ReplaceM, ReplaceO, ..), where we also added a formal definition of what those tags describe in terms of linguistic errors.
+We created the repository using the error tags, where we also added a formal definition of what those tags describe in terms of linguistic errors.
 In addition, we expanded the error taxonomy by splitting it into finer-grained classes.
@@
 (2013), which was refined later and introduced as ARETA Belkebir and Habash (2021).
+Appendix C presents examples of the error types.
 Figure 2 shows some examples from our error instruction repository.
@@
 The model processed one essay at a time in a zero-shot setting, except that we included the definition and explanation of the error tags.
-For example, Replace_M indicates a morphological error, while Merge targets two mistakenly split tokens that need to be merged, and so on.
+For example, M indicates a morphological error, while Merge targets two mistakenly split tokens that need to be merged, and so on.
 After conducting multiple experiments, we observed the following issues: (i) The model struggled to follow the predefined error distribution perhaps due to the complexity of the prompts.
-(ii) The model was confused by certain error tags, particularly SPLIT, MERGE as well as the UNK (unknown errors).
+(ii) The model was confused by certain error tags, particularly Split and Merge.
 These errors were mainly ignored in the injected text.
 (iii) We calculated the cosine similarity between the main error profile and the injected essays’ error distributions as shown in Equation 3.
-When we injected all errors at once, the similarity agreements did not exceed 20%; however, when we reduced the number of error tags per essay or omitted the misleading tags, the agreements significantly improved, reaching 86% Therefore we implemented a method where each error type was injected separately.
+When we injected all errors at once, the similarity agreements did not exceed 20%; however, when we reduced the number of error tags per essay the agreements significantly improved, reaching 86%.
+Therefore we implemented a method where each error type was injected separately.
 This required multiple iterations over the same essay, corresponding to the number of error tags shown in the error distribution profile for each CEFR level.
 Figure 3 shows an example of a GPT-4o prompt for error injection.
-Some errors occurred more frequently than others among Arabic writers.
-For example, orthographic errors are more frequent than semantic errors.
-So, rather than repeating a single error instruction for all instances of an error tag, we randomly select a weighted number of error instructions proportional to the average occurrence of each error type.
+Some error types, especially orthographic errors, are more frequent among Arabic writers than others.
+The prompt was intentionally designed through prompt engineering.
+The ‘helpful assistant’ component establishes a cooperative persona for the LLM, while the subsequent instruction to ‘inject erroneous tokens’ explicitly guide GPT-4o towards the specific task of error introduction.
+This approach ensures that GPT-4o is not making random edits but is rather following predefined instructions to create targeted errors, aligning with the overall goal of generating realistic synthetic data.
+To reflect this, we randomly select weighted error instructions based on the average frequency of each error type.
 Figure 4 shows the pseudocode for the selection process.
-The full pseudocode is in Appendix A.
+The full pseudocode is in Appendix B.
 Figure 3: Sample GPT-4o error injection prompt
@@
 Train Data QWK Acc F1 P R QWK Acc F1 P R
-Zaebuc-1 22.44 57.58 24.50 23.33 26.76 61.06 84.85 43.70 42.50 45.31
-Zaebuc-1 + GPT essays 14.92 60.61 26.43 25.55 27.45 96.00 96.97 92.32 98.04 88.89
-Zaebuc-1 + BERT errors 27.87 57.58 38.02 35.86 44.93 82.70 87.88 71.66 70.83 74.38
-Zaebuc-1 + GPT errors_1 17.14 57.58 25.64 25.18 26.27 96.47 96.97 94.16 97.92 91.67
-Zaebuc-1 + GPT errors_0 20.84 57.58 32.76 31.53 46.08 93.79 93.94 95.12 96.49 94.44
+ZAEBUC (baseline) 22.44 57.58 24.50 23.33 26.76 61.06 84.85 43.70 42.50 45.31
+ZAEBUC + GPT essays 14.92 60.61 26.43 25.55 27.45 96.00 96.97 92.32 98.04 88.89
+ZAEBUC + BERT errors 27.87 57.58 38.02 35.86 44.93 82.70 87.88 71.66 70.83 74.38
+ZAEBUC + GPT errors_1 17.14 57.58 25.64 25.18 26.27 96.47 96.97 94.16 97.92 91.67
+ZAEBUC + GPT errors_0 20.84 57.58 32.76 31.53 46.08 93.79 93.94 95.12 96.49 94.44
 Table 6: Performance comparison of different training datasets.
-GPT essays are the original correct essays generated from GPT-4o, BERT errors are the erroneous essays using the controlled injection BERT model, GPT errors_1 are the erroneous essays using GPT-4o with one shot error example, while GPT errors_0 with Zero shot settings
-6 Evaluation
-6.1 Metrics
-Our primary evaluation metric is Quadratic Weighted Kappa (QWK) Cohen (1968), the most widely-used metric in recent AES research Ke and Ng (2019).
+GPT essays are the original correct essays generated from GPT-4o, BERT errors are the erroneous essays using the controlled injection BERT model, GPT errors_1 are the erroneous essays using GPT-4o with one-shot error example, while GPT errors_0 with Zero-shot settings.
+6 Experimental Setup
+This study focuses on introducing a data augmentation framework and synthetic Arabic essay corpus, rather than proposing a new AES model.
+We use a BERT-based model trained on the original ZAEBUC dataset as the reference baseline, evaluating how different augmentation strategies (e.g., GPT-4o generation, BERT-based error injection) improve performance relative to this setup.
+6.1 Data and Metrics
+We use the ZAEBUC dataset for all the experiments, following the splits created by Alhafni et al.
+(2023): 70% Train, 15% Dev, and 15% Test.
+Our primary evaluation metric is Quadratic Weighted Kappa (QWK) Cohen (1968), the most widely used metric in AES research Ke and Ng (2019).
 We also report accuracy, macro precision (P), recall (R), and F1 scores.
-Model predictions are evaluated against the average of the gold holistic CEFR labels and in a multi-reference setting against the three individual gold labels per essay (§3).
-6.2 Data
-We use the ZAEBUC dataset for all the experiments.
-The dataset was split (70%,15%,15%) for Train, Dev and Test sets, respectively Alhafni et al.
-(2023).
-Table 6 shows the performance results of the BERT-based Scoring Model utilizing different training sets.
+Model predictions are evaluated in two settings settings: average-refence and multi-reference.
+The average-reference setting uses the rounded average of the three scores as the gold label, while the multi-reference considers each of the three human-assigned labels as a valid reference during evaluation, following a more tolerant evaluation strategy (§3).
 6.2 Model
-We treat the task of automated essay scoring (AES) as a text classification problem.
-For Arabic AES, we fine-tune CAMeLBERT MSA Inoue et al.
+We treat the task of AES as a text classification problem.
+We fine-tune CAMeLBERT MSA Inoue et al.
 (2021) on the training split of ZAEBUC.
@@
 Our baseline system, only trained on the ZAEBUC training set, indicates room for improvement, with the F1 at 24.50% and QWK at 22.44%.
-We then switched between different datasets to enhance the Arabic AES model and to measure the impact of data augmentation on the model.
-We introduced the 3,040 correct essays generated using GPT-4o, which dropped the QWK compared to the baseline while increasing the F1.
-However, we observed significant improvements in the multi-reference settings for the QWK (96.00%) and F1 (92.32%).
-Since this is an error-free, synthetic dataset, we further refine the model by introducing essays that mimic human-like writing styles.
-We compare the two approaches introduced in section 5: (1) GPT-based error injection (with and without error instruction examples) and (2) the controlled error injection method based on CAMeLBERT MSA.
-Inoue et al.
-(2021).
-The results demonstrate that the controlled error model enhances performance across all metrics, particularly in the average-reference setting, that achieved 27.87 % and 38.02 for QWK and F1 respectively.
-However, the GPT-based error injection performs best under the multi-reference setting, where the one-shot essays increased the QWK to 96.47 and the zero-shot erroneous essays improved the F1 to 95.12%.
-These results suggest that LLMs and Transformer methods generalize effectively in AES models and also improve performance.
-The results emphasize that a controlled methodology is more reliable in mimicking patterns of learner errors and mistakes.
-This positions the controlled BERT-based model as a vital component of the data augmentation pipeline, effectively addressing the limitations of the GPT models.
+We then switched between different datasets to measure the impact of data augmentation on the model.
+Impact of Synthetic Data
+We tested data augmentation by adding 3,040 corrected GPT-4o-generated essays, which lowered QWK but increased F1.
+Notably, the multi-reference setting saw significant gains, with QWK at 96.00% and F1 at 92.32%.
+This pattern stems from the flexibility of multi-reference evaluation, which treats all three human-assigned CEFR labels as valid references.
+This accommodates natural scoring variations and increases the chance that model predictions, especially on synthetic data, align with at least one reference label, boosting QWK and F1 scores for both GPT-generated and error-injected essays.
+Comparison of Error Injection Methods
+As the initial synthetic essays were error-free, we further refined the model by adding essays with human-like errors.
+We compared two methods from §5: (1) GPT-based error injection (with and without instruction examples) and (2) the controlled BERT-based method.
+The results demonstrate that the controlled error model improves performance in all metrics, particularly in the average reference setting, which achieved 27.87 % and 38.02 for QWK and F1, respectively.
+This result aligns with expectations, as the BERT-injected errors closely follow CEFR-based error distributions, producing errors that realistically reflect learner writing and better match the average of human ratings.
+GPT-based error injection performed best in the multi-reference setting, with one-shot examples reaching 96.47% QWK and zero-shot boosting F1 to 95.12%.
+While less aligned with CEFR profiles, GPT errors benefit from fluency and variability, increasing the chance of matching at least one human reference in this flexible evaluation.
+7 Discussion
+This study demonstrated the effectiveness of synthetic data and controlled error injection in enhancing Arabic AES, providing key insights into metric interpretation, data expansion, and methodological choices.
+First, we emphasize that QWK offers a more robust metric than accuracy for evaluating AES systems, particularly under imbalanced class distributions.
+Unlike accuracy, which is biased by majority classes, QWK penalizes errors by their ordinal distance from the correct label.
+As Table 6 shows, even modest QWK improvements indicate meaningful advancements in differentiating CEFR levels, a distinction especially relevant given the skewed ZAEBUC dataset.
+The significant gains observed in the multi-reference setting with generated GPT-4o essays stem from its flexibility.
+This evaluation approach treats all three human-assigned CEFR labels as valid references, accommodating natural scoring variations and increasing the chance that model predictions align with at least one reference label.
+Our analysis revealed that while GPT-4o is powerful for generating diverse content, it struggles to precisely follow the nuanced distribution and specific linguistic features, including error patterns, observed in the manually annotated ZAEBUC dataset.
+In the GPT-based error injection approach, error type selection is guided by average error counts from the ZAEBUC corpus, but error realization depends on GPT-4o’s interpretation of the prompt, making it less predictable.
+This inherent challenge in mimicking human-like linguistic and error distributions through zero-shot generation directly contributed to the observed lower agreement rate.
+In contrast, the controlled method employs a BERT-based classifier for error prediction and applies transformations using bigram-MLE.
+This systematic approach resulted in a more robust replication of empirically observed error patterns, leading to its superior performance in the average-reference setting.
+This is expected, as BERT-injected errors more closely resemble learner writing and align more closely with average human ratings.
+Overall, our findings highlight a trade-off between error alignment and fluency in data augmentation.
+Controlled error injection excels in the average-reference setting due to its closer alignment with learner errors, while GPT-based augmentation benefits from multi-reference flexibility but less reliably replicates authentic errors.
+The controlled BERT-based method thus serves as a key component of our pipeline, effectively addressing the limitations of direct GPT error injection.
+Qualitative Analysis
+The qualitative analysis of the generation process revealed various biases in the GPT-4o outputs, including cultural, gender, and ideological biases.
+For instance, the essays frequently referenced traditional Arabic themes, reinforced stereotypical gender roles, and reflected culturally narrow assumptions.
+A clear example of religious bias is that \<الجمعة> ‘Friday’ was selected as the favorite day in all 20 generated essays.
+Additionally, there was a noticeable tendency to use masculine forms throughout the texts.
+Such biases may unintentionally disadvantage students whose writing reflects different experiences, perspectives, or identities.
+Examples of these biases, along with their frequencies, are provided in Appendix A.2.
+We also observed a lack of diversity among the ten essays generated per prompt, with GPT-4o often repeating similar lexical and structural patterns.
 8 Conclusions and Future Work
-This paper proposes a hybrid framework for building an Arabic AES system.
-We utilize both LLMs and Transformers to overcome the challenges of dataset scarcity by generating synthetic data that mimics the writing styles of Arabic learners.
-We leverage the ZAEBUC corpus and build linguistic feature profiles as well as error profiles that align with each CEFR score.
-We prompt GPT-4o to generate 3,040 Arabic essays based on multiple CEFR levels corresponding to 152 prompts.
-However, GPT-4o’s effectiveness depends heavily on prompt engineering, achieving only 27.5% agreement with the original linguistic feature profiles.
-To introduce errors into GPT-generated essays, we compare two approaches: (1) GPT-4o error injection, which struggles with realistic learner writing due to misinterpreted linguistic errors or complex prompts, causing inconsistencies; (2) a controlled method fine-tuning the CAMeLBERT MSA model to inject errors proportionally to their natural occurrence.
-The latter method is an essential component of data augmentation for Arabic AES.
-Our evaluation, using a fine-tuned BERT classifier, shows that the proposed hybrid framework that combines GPT-generated data and a controlled error injection method shows consistent improvement over the baseline (QWK: 27.87%, F1: 38.02%), achieving more predictable and interpretable results.
-These results suggest the controlled approach to generating learner errors better captures the actual distribution of errors across proficiency levels.
-In future work, we plan to expand the dataset by incorporating real essay samples, which will enhance the linguistic and error profiles for each CEFR level, thereby improving the proposed data augmentation model.
-In addition, we will deploy the AES system as part of an interactive educational writing assistant tool to help students enhance their Arabic writing with immediate feedback on errors and CEFR scores.
-8 Limitations
+This paper presents a hybrid framework for Arabic AES, using LLMs and transformers to tackle data scarcity by generating synthetic essays that partly replicate Arabic learner writing.
+Building on the ZAEBUC corpus, we developed CEFR-aligned linguistic and error profiles and used GPT-4o to produce 3,040 essays across 152 prompts.
+However, GPT-4o’s performance relies heavily on prompt engineering, achieving only 27.5% alignment with our reference profiles.
+To introduce errors, we compare our two methods: (1) GPT-4o prompted multi-step error injection, and (2) our controlled method fine-tuning the CAMeLBERT MSA model to inject errors proportionally to their profiled occurrence.
+Evaluated with a fine-tuned BERT classifier, our hybrid framework, combining GPT-generated data with controlled error injection, outperformed the baseline (QWK: 27.87%, F1: 38.02% ), offering more reliable and interpretable results.
+These findings demonstrate the effectiveness of controlled error injection in capturing learner error distributions across CEFR levels.
+For future work, we will prioritize integrating a human evaluation into our framework.
+Human annotators will assess the fluency and naturalness of synthetic essays, as well as the realism of injected errors, ensuring that they reflect typical learner patterns at specific CEFR levels.
+To improve generalizability, we also plan to expand the diversity of prompts beyond predefined topics and incorporate a wider set of writing traits, including coherence, logical flow, and topic relevance, beyond syntactic and lexical features.
+We also intend to enhance CEFR-level modelling by incorporating more manually annotated essays.
+This will help capture nuanced linguistic variations across levels and increase the robustness of our dataset.
+Lastly, we aim to deploy the AES system as an interactive tool to provide users with instant feedback on errors and proficiency levels.
+Limitations
 Despite the effectiveness of our hybrid Arabic AES framework, we note several limitations related to the quality of generated Arabic essays, error injection accuracy, and the generalization of the AES model.
 The lack of A1 and C2 essays in ZAEBUC means that there is no gold reference data for these levels, which may impact both linguistic and error profiles, affecting the accuracy of GPT-generated essays.
-Furthermore, different biases exist in both ZAEBUC and the GPT outputs: ZAEBUC overrepresents B-level essays, skewing model training; and GPT-4o tends to favor Arabic culture themes, thereby limiting diversity in the generated texts, and usability in other contexts.
-In addition, due to the lack of comprehensive gold data, GPT struggles to fully replicate real learner writing styles, achieving only 27.5% agreement with linguistic feature profiles.
+Furthermore, different biases are present in both the ZAEBUC dataset and GPT-4o outputs as discussed in (§7)In addition, due to the lack of comprehensive gold data, GPT struggles to fully replicate real learner writing styles, achieving only 27.5% agreement with linguistic feature profiles.
 Another limitation is the model’s ability to generalize across various domains and question types.
@@
 Moreover, better-controlled GPT prompting could refine the quality and diversity of generated essays, reducing biases and improving alignment with real learner writing patterns.
-9 Ethical Considerations
+Due to resource constraints, human evaluation was not feasible in this study; however, we plan to engage CEFR-trained annotators in the future.
+Ethical Considerations
 While Arabic AES systems provide significant support in assessing Arabic learners’ writing proficiency, it is essential to highlight the ethical implications of their use.
@@
 AES tools should serve as an educational assistive technology, complementing the teacher’s judgment, not replacing it in educational settings.
-Appendix A GPT-4o Error injection algorithm
-Figure 6: GPT-4o Error injection and verification algorithm
+Appendix A GPT-4o Generated Essays
+A.1 Statistics
+CEFR #Essays #Words #Sentences #Tokens Avg_W_L Avg_S_L Unique_Tokens Unique_Words
+A1 470 39367 6721 47612 4.66 6.08 4518 4512
+A2 470 54980 5565 63769 4.82 10.46 7147 7143
+B1 530 100185 7276 113672 4.97 14.62 10742 10734
+B2 530 127029 7804 142995 5.03 17.32 12522 12509
+C1 520 136849 7630 152962 5.16 19.05 12832 12823
+C2 520 146253 7946 163461 5.16 19.57 13686 13676
+Total 3040 604663 42942 684471 5.04 14.94 27286 27272
+Table 7: Summary statistics of the generated Arabic synthetic essay corpus across CEFR levels.
+#Essays denotes the number of essays; #Words refers to the total word count; #Sentences indicates the total number of sentences; #Tokens represents the total number of tokens (vocabulary items); Avg_W_L corresponds to the average word length in characters; Avg_S_L refers to the average sentence length in words; lexical diversity is captured through the counts of unique tokens and unique words.
+A.2 Bias in the Generated Essays
+Arabic Prompt English Prompt Arabic Response English Response Occurrences /20 Bias
+\< ما هو يومك المفضل؟ > What is your favorite day? \<الجمعة > Friday 20 Cultural/Religious Bias
+\< ماهو طعامك المفضل؟ > What is your favorite food? \< البيتزا > Pizza 16 Globalization Bias
+\< ما هي هوايتك المفضلة؟ > What is your favorite hobby? \< القراءة > Reading 10 Socioeconomic/Class Bias
+\< رحلة ذهبت إليها > A trip you went on \< ذهبت إلى البحر > I went to the beach 18 Geographical/Cultural Bias
+\< ماذا تفعل في عطلة نهاية الأسبوع؟ > What do you do on the weekend? \< نذهب إلى الحديقة > We go to the park 20 Geographical/Cultural Bias
+\< ماهي المادة الدراسية المفضلة؟ > What is your favorite school subject? \< الرياضيات > Mathematics 17 Educational System Bias
+\< ما هي رياضتك المفضلة؟ > What is your favorite sport? \< كرة القدم > Football 20 Cultural Bias
+\< شخص تعتبره مثلك الأعلى > A person you consider your role model \< والدي > My father 17 Gender Bias
+\< من هو أفضل صديق لك؟ > Who is your best friend? \< أحمد > Ahmed 20 Gender/Name Bias
+\< المهنة المستقبلية > Your future profession \< طبيب > Doctor 13 Stereotype Bias
+\< معلم شهير في العالم الأدبي > A famous teacher in the world of literature \< الأهرامات > The Pyramids 17 Cultural Bias
+\< أفضل فصول السنة > Your favorite season of the year \< الصيف > Summer 20 Climate Bias
+\< لغة تود تعلمها > A language you would like to learn \<الأسبانية > Spanish 16 Language Bias
+\< بلد ترغب في السفر إليه > A country you would like to visit \< مصر > Egypt 10 National Identity Bias
+\< قوة خارقة تتمناها > A superpower you wish to have \< الطيران > Flying 19 Media Bias
+Table 8: Examples of response biases in GPT-4o generated essays.
+Appendix B GPT-4o Error Injection Algorithm
+Appendix C Error Types Taxonomy
+11-Classes 42-Classes Error Description Correct Word Erroneous Word
+Morphology (M) M MI Inflection \< عارف> \< معروف>
+MT Tense \< ذهب> \< يذهب>
+Orthography (O) O OA Alef-Maqsura \< القاضي> \< القاضى >
+OA+OH Alef-Maqsura + Hamza \< أضحى > \< اضحا >
+OA+OR Alef-Maqsura + Wrong Character \< كشيء> \< كشىء>
+OC Chatacter Order \< المدرسة > \< المردسة>
+OD Extra Character \< هذا> \< هاذا>
+OD+OG Extra Character + Lengthening Short Vowels \< تتطورو > \< تتطور >
+OD+OH Extra Character + Hamza \< لأنهم > \< ألانهم >
+OD+OM Extra Character + Missing Character \< الاجتماعي > \< الاجتاعيي >
+OD+OR Extra Character + Wrong Character \< الصور > \< السوور>
+OH Hamza \< العب > \< إلعب >
+OH+OM Hamza + Missing Character \< الأشياء > \< الاشيا >
+OH+OT Hamza + Ta-Marbuta \< إمارة > \< اماره >
+OM Missing Character \< المدرسة > \< المدسة >
+OM+OR Missing Character + Wrong Character \< المجتمع > \< الجطمع >
+OR Wrong Character \< المدرسة> \< المدرصة >
+OR+OT Wrong Character + Ta-Marbuta \<مكتظة> \< مكتضه>
+OT Ta-Marbuta \< غرفة> \< غرفه >
+OW Alef-Fariqa \< كتبوا > \< كتبو >
+Semantics (S) S SF Conjunction \< فسبحان > \< سبحان >
+SW Word Selection \< على > \< من >
+Punctuation
+P P Punctuation \<السوق،> \<السوق.>
+Syntax (X) X XC Case \<رائعا> \<رائع>
+XC+XG Case + Gender \<مجتهدا> \<مجتهدة>
+XC+XN Case + Number \<نواح> \<نواحي>
+XF Definiteness \<المفيد> \<مفيد>
+XG Gender \<كان> \<كانت>
+XM Missing Word \<على> NULL
+XN Number \<كتابين> \<كتب>
+XT Unnecessary Word NULL \<على>
+Combination (Comb.) M+O MI+OH Inflection + Hamza \< أشخاص > \< اشخاصك >
+O+X OH+XC Hamza + Case \< أضرارا> \< اضرار >
+SPLIT SPLIT SPLIT Split \<دولة الإمارات> \<دولةالإمارات>
+MERGE MERGE MERGE Merge \< بالعلم > \< ب العلم >
+DELETE DELETE DELETE Delete NULL \<داخل>
+Table 9: Illustrative examples of error types categorized according to the ARETA error taxonomy Belkebir and Habash (2021).
+The table presents hierarchical mappings from coarse-grained (11-Class) to fine-grained (42-Class) error categories, alongside representative corrections.